Calculate the pH of a solution of 0.25 M RbBrO2. Ka HBrO2 = 1.2 x 10-5
1
Expert's answer
2020-07-24T14:52:04-0400

Answer on Question #127359, Chemistry, General Chemistry


Calculate the pH of a solution of 0.25 M RbBrO2. Ka HBrO2 = 1.2 x 10-5


Solution:


RbBrO2 is a salt which consists of cation of the strong base RbOH and anion of the weak acid HBrO2. Hydrolysis reaction is:

RbBrO2 + H2O = RbOH + HBrO2 - molecular equation

BrO2- + H2O = OH- + HBrO2 - ionic equation (pH > 7, basic medium)

The pH value of this salt can be calculate as:

pH = 7 + (рКа - рС)/2, where:

pKa = - lg Ka = - lg (1.2 x 10-5) = 4.92

pC = - lg C(RbBrO2) = - lg 0.25 = 0.60

So:

pH = 7 + (4.92 - 0.60)/2 = 9.16


Answer: pH = 9.16.